# From Canopy to Ground via ForestGen3D: Learning Cross-Domain Generation of 3D Forest Structure from Aerial-to-Terrestrial LiDAR

### v1

@@ 3 Experimentation
The experiments in this section are designed to address two key questions: (1) Can conditional DDPM models effectively learn the detailed 3D structural distribution of forest vegetation from TLS data with ALS guidance? and (2) Is the approach scalable to plot-scale and landscape-scale deployment, enabling accurate synthesis of tree-level structure and improvement of downstream ecological biometrics commonly used in forest management?. The first set of experiments, presented in Section 3.1, evaluates the generative capacity of DDPMs at the tree scale. We assess whether the model’s inductive biases and architecture are well-suited for representing the hierarchical and self-similar structure of vegetation. This includes analysis of the model’s approximation quality via reconstruction error, as well as ablation studies that explore the sensitivity of generation quality to key DDPM hyperparameters such as diffusion steps.To quantitatively assess reconstruction error, we compute Chamfer Distance (CD) and Earth Mover’s Distance (EMD) between generated and reference TLS point clouds, using held-out tree samples. In addition, we examine the relationship between CD, EMD, and our expected point containment (EPC) metric . The goal of this analysis is to evaluate whether containment can serve as a practical proxy for CD and/or EMD in scenarios where ground-truth TLS data is unavailable. The second set of experiments, detailed in Sections 3.3-3.5, demonstrates the model’s generalization and scalability from tree to landscape scale scenarios in mixed conifer ecosystems. At the tree and plot scales, we evaluate the consistency of ALS-guided generation in matching spatial extents and density patterns observed in ground truth TLS scans. At the regional scale, we deploy the model across 200 meter radius ALS-only regions and assess the realism of synthesized vegetation. For both plot and landscape cases, we compute ecological biometrics including tree height, diameter at breast height (DBH), crown diameter (CrD), and crown volume (CrV) based on ALS, TLS, ALS+TLS, and ALS+ForestGen3D samples. These comparisons highlight the model’s ability to augment sparse ALS with realistic sub-canopy structure, leading to improved and more complete biometric estimates. Furthermore, we confirm that the containment prior is preserved in inference-time landscape-scale deployments, as evidenced by low out-of-envelope deviation statistics across diverse mixed-conifer landscapes.
@@ 3.1 Training and Validation
To evaluate the inductive biases and approximation fidelity of our DDPM-based generative framework, we conduct a series of ablation experiments. During training, each batch consists of 16 randomly selected ALS/TLS example pairs. Each point cloud is uniformly subsampled to a fixed size of $N=2048$. This random subsampling enhances dataset diversity by introducing spatial variability across training epochs, enabling the model to encounter a broader range of structural configurations over time. Although the $N=2048$ resolution may appear limiting, it applies only at the level of individual trees. When aggregated over larger spatial extents such as plots or landscapes, this per-tree resolution naturally scales to yield detailed reconstructions across entire forested regions. Moreover, higher-resolution representations can be achieved through ensemble sampling, in which multiple stochastic generations are produced for the same conditional ALS input and merged coherently to create denser 3D tree structures. As an illustration of the capabilities of the generative model to approximate 3D tree structures given the existing settings, we include Figure 6 showcasing three examples in the training set with ALS input (first column), AI-generation (second column) and ground truth TLS (third column) in each example. Here, we see that our model is able to generate samples at good qualitative approximations of the 3D structural features of TLS distributions conditional on the ALS inputs across the entire vertical tree-stand direction. Figures 6(a) and 6(b) depict instances where the ALS input contains a single tree, which the model accurately approximates, demonstrating its ability to infer complex internal structures from sparse input. Figure 6(c) further illustrates its performance in closely packed tree settings. Here, ground truth TLS shows four trees with tightly packed canopies, while the model generates fewer distinct trunks than the ground truth. Despite this, the output remains visually realistic and spatially consistent with what a human observer might infer from the incomplete ALS scan.
@@
Figure 6: Three examples of DDPM generation at the tree-scale conditioned by ALS inputs in the training set. This qualitative examples showcase the architecture’s capacity to encode and represent the 3D structural details of trees.
Ablation studies to analyze model reconstruction performance where conducted by varying the number of diffusion steps $T\in\{$200, 400, 600, 1000, 2000}\}, diffusion parameters $\beta_{0}=10^{-4}$, $\beta_{T}=0.02$, and batch size 16 constant. These experiments aim to analyze the error of the model at training time and also the spatial containment consistency in eq.(2). We track reconstruction error across training iterations using both Chamfer Distance (CD) and Earth Mover’s Distance (EMD), computed between generated and ground-truth TLS point clouds on a held-out validation set of 300 tree examples. As shown in Figs. 7(a)-7(b), both CD and EMD consistently decrease with training iterations across all diffusion schedules, indicating convergence and improving fidelity. Notably, models with higher diffusion steps (e.g., $T=1000$) exhibit slower convergence but ultimately achieve slightly lower reconstruction error, while smaller $T$ values converge quickly with slightly higher error floors.
@@
Figure 7: Validation set reconstruction error as a function training iterations.
@@
(a) Expected Point Containment Consistency (EPC)
(b) Chamfer Distance versus EPC
(c) Earth Movers Distance versus EPC
Figure 8: Validation set EPC versus training iteration.
@@ 3.2 Testing dataset
For testing, we used 50 TLS scans collected from geographic regions entirely outside the training and validation areas, though sharing similar ecosystem characteristics and collected with the same sensors used for the training data. We evaluated the model’s performance across three spatial scales. At the tree scale, we followed the same co-registration and tree detection procedures used during training, resulting in 1,457 individual test trees. This setup enabled direct, one-to-one comparisons between generated samples and corresponding TLS ground truth reference data at high structural fidelity. At the plot scale, we constructed test plots by aggregating ALS data over spatial extents corresponding to each TLS scan, truncating ALS coverage to match the TLS field of view (e.g., 25 meters in radius ). This allowed for evaluation of the model’s ability to reconstruct localized forest structure. Finally, at the regional scale, we assessed scalability and extrapolation by deploying the model over multiple 200 meter radius areas containing ALS data. These larger, fully independent areas tested the model’s ability to synthesize realistic vegetation structure across broader, ALS only landscapes. This evaluation strategy enables comprehensive assessment of ForestGen3D’s robustness and flexibility from fine-grained, tree-level detail to broad-scale generation in mixed conifer ecosystems.
@@ 3.3 Tree-Scale ALS-Guided Generation
In Fig. 9 we include three representative qualitative comparisons from the test set to illustrate differences between completion/generation approaches across a range of tree structures. First two columns across all three cases represent ALS input and TLS ground truth, while the remaining columns correspond to the completion/generation results. In all three examples, PCN tends to oversmooth crowns, recovering coarse silhouettes while missing high-frequency branching detail. The other baseline generative models (3D-GAN, latent-GAN, and PointFlow) produce shapes that appear closer to the TLS shape, yet are less consistent across examples in general, in comparison to those from our ForestGen3D model. For instance, in the second example, latent-GAN (l) yields a form whose stem does not resemble the straight trunk structure of the TLS reference. Similarly, PointFlow in (f) generates a discontinuous tree stem of larger a diameter compared to ground truth TLS. 3D-GAN and ForestGen3D produce consistently closer results to TLS with canopy structure close to that of ALS and coherent stem-to-crown transitions, recognizable branching patterns, and overall shapes that better match human expectations of realistic tree architecture. These qualitative outcomes mirror the quantitative results below each figure, highlighting the capabilities of generative models to produce realistic structural properties of forests.
@@
Figure 9: Qualitative comparison of ALS-conditioned tree structure generation across three representative examples. (a,h,o) ALS input and (b,i,p) TLS reference; (c,j,q) PCN [14], (d,k,r) 3D-GAN [15], (e,l,s) latent-GAN [16], (f,m,t) PointFlow [17], and (g,n,u) ForestGen3D (ours). PCN oversmooths crowns, producing coarse silhouettes with little fine-scale detail. The other baselines generate shapes that more closely resemble TLS references but with inconsistencies such as distorted trunks (latent-GAN in (l)) or more discontinuous stems (PointFlow in (f)). 3D-GAN and ForestGen3D yield more coherent reconstructions, with ForestGen3D showing smaller metric errors with respect to ground.
@@
Table 1: Generation/completion errors in test set.
@@
Table 1 shows a quantitative evaluation summarizing the CD/EMD and EPC metrics over the test dataset of 1457 test examples. This table summarizes the resulting metrics for the completion-based PCN [14] method and the generative models including 3D-GAN [15], latent l-GAN [16], PointFlow [17] and our ForestGen3D denoising diffusion based model. For reference, example CD/EMD/EPC values are also reported in Fig. 9 for their corresponding examples. Beyond the numerical values, the metrics highlight important differences between the approaches. PCN, while competitive in CD, performs worst in EMD, reflecting its bias toward capturing coarse, low-frequency geometry at the expense of fine structural detail. Latent-GAN shows the largest CD, due to plausible over-compressing structural features in the latent layers, while 3D-GAN improves CD, it still remains above our model. PointFlow achieves stronger balance across CD and EMD, though it is also higher than the performance of our model. In contrast, ForestGen3D consistently achieves the lowest CD and EMD, while also attaining the highest EPC, indicating superior alignment with TLS reference structure and better recovery of fine-scale details. These results demonstrate that ForestGen3D more effectively learns the underlying distribution of forest structure from ALS–TLS data, achieving overall improved performance compared to both completion/generation-based baselines.
@@ 3.4 Plot-scale ALS-Guided Generation
We also conduct both qualitative and quantitative analyses at the plot scale, using circular areas with a 25 meter radius. For the qualitative evaluation, we visually compare structural reconstructions obtained from ALS, TLS, ALS+TLS and a hybrid approach combining ALS with ForestGen3D generation enhancement. For each plot, individual trees were first detected using 3D bounding boxes, and our ForestGen3D model was then applied to generate plausible understory structure beneath each detected tree. Regions lacking tree or shrub like vegetation are left unaltered under the assumption that they do not suffer from occlusion in ALS. To support the quantitative assessment, we perform an evaluation of vegetation biometric estimates derived from each data source: ALS alone, TLS alone, ALS+TLS and ALS enhanced with ForestGen3D. The combined ALS+TLS data serves as ground truth, given its ability to capture detailed, high-resolution vegetation structure across the entire vertical extent of the tree, including the ground, overstory and understory components.
Figure 10 illustrates four representative testset examples (one per row) at plot-scale. The first column illustrates the ALS point cloud, the second column shows the corresponding co-registered ground truth TLS scan, and the third column presents the results of aggregating the ALS scan with the generated under-story from our ForestGen3D model. Note that in all examples, the ALS plots suffer from either missing or highly sparse understory information between the ground and tree foliage, as can be observed in the images in the first column where points are missing. TLS, on the other hand, provides high resolution in the elevation/altitude direction, as shown in the second column, where the main tree stem is finely resolved along the vertical tree-stand direction. The third column, showing the point cloud from ALS+ForestGen3D, illustrates that our approach is capable of enhancing or generating realistic understory structure at the plot scale, where the generated stems and ground features fill in gaps and transition smoothly from the available ALS data.
@@
(l) Ex.4:ALS + ForestGen3D.
Figure 10: Representative examples comparing 3D structure from ALS, TLS and ALS +ForestGen3D.

### v2

@@ 3 Experimentation
The experiments in this section are designed to address two key questions: (1) Can conditional DDPM models effectively learn the detailed 3D structural distribution of forest vegetation from TLS data with ALS guidance? and (2) Is the approach scalable to plot-scale and landscape-scale deployment, enabling accurate synthesis of tree-level structure and improvement of downstream ecological biometrics commonly used in forest management?. The first set of experiments, presented in Section 3.2, evaluates the generative capacity of DDPMs at the tree scale. We assess whether the model’s inductive biases and architecture are well-suited for representing the hierarchical and self-similar structure of vegetation. This includes analysis of the model’s approximation quality via reconstruction error, as well as ablation studies that explore the sensitivity of generation quality to key DDPM hyperparameters such as diffusion steps. To quantitatively assess reconstruction error, we compute Chamfer Distance (CD) and Earth Mover’s Distance (EMD) [33] between generated and reference TLS point clouds, using held-out tree samples. In addition, we examine the relationship between CD, EMD, and our expected point containment (EPC) metric . The goal of this analysis is to evaluate whether containment can serve as a practical proxy for CD and/or EMD in scenarios where ground-truth TLS data is unavailable. The second set of experiments, detailed in Sections 3.4-3.6, demonstrates the model’s generalization and scalability from tree to landscape scale scenarios in mixed conifer ecosystems. At the tree and plot scales, we evaluate the consistency of ALS-guided generation in matching spatial extents and density patterns observed in ground truth TLS scans. At the regional scale, we deploy the model across 200 meter radius ALS-only regions and assess the realism of synthesized vegetation. For both plot and landscape cases, we compute downstream ecological biometrics including tree height, diameter at breast height (DBH), crown diameter (CrD), and crown volume (CrV) based on ALS, TLS, ALS+TLS, and ALS+ForestGen3D samples. These comparisons highlight the model’s ability to augment sparse ALS with realistic sub-canopy structure, leading to improved and more complete biometric estimates. Furthermore, we confirm that the containment prior is preserved in inference time landscape scale deployments, as evidenced by low out of envelope deviation statistics across diverse mixed-conifer landscapes.
3.1 Implementation Details
ForestGen3D is implemented in Python/PyTorch with CUDA acceleration. All experiments were conducted on an NVIDIA Tesla V100S-PCIE-32GB GPU, with a total training time of approximately 20 hours for 100 epochs. The diffusion process follows a standard DDPM formulation, using a linear noise schedule with $\beta_{0}=1\times 10^{-4}$, $\beta_{T}=2\times 10^{-2}$, and $T=1000$ diffusion steps, although we also test other $T$ settings. The model is trained using a mean squared error (MSE) loss, while the Earth Mover’s Distance (EMD) and Chamfer distance metric is optionally monitored during validation to ensure that the learned denoising process maintains geometric stability. Absolute coordinates, local neighborhood features, global shape descriptors, and ALS conditioning are included throughout the denoising process. Grouped feature attention, batch normalization, and residual connections are enabled across all modules to stabilize training. Radius neighborhood search is used throughout both encoding and decoding, while a $k$-nearest neighbors (kNN) strategy with $k=8$ is employed for feature propagation during decoding. At inference, this configuration produces a complete 3D tree structure in approximately 0.4 seconds. To support reproducibility and reuse, the ForestGen3D model implementation, trained weights, and associated configuration files are expected to be released publicly upon publication of the manuscript.
@@ 3.2 Training and Validation
To evaluate the inductive biases and approximation fidelity of our DDPM-based generative framework, we conduct a series of ablation experiments. During training, each batch consists of 16 randomly selected ALS/TLS example pairs. Each point cloud is uniformly subsampled to a fixed size of $N=2048$. This random subsampling enhances dataset diversity by introducing spatial variability across training epochs, enabling the model to encounter a broader range of structural configurations over time. Although the $N=2048$ resolution may appear limiting, it applies only at the level of individual trees. When aggregated over larger spatial extents such as plots or landscapes, this per tree resolution naturally scales to yield detailed reconstructions across entire forested regions. Moreover, higher-resolution representations can be achieved through ensemble sampling, in which multiple stochastic generations are produced for the same conditional ALS input and merged coherently to create denser 3D tree structures. As an illustration of the capabilities of the generative model to approximate 3D tree structures given the existing settings, we include Figure 6 showcasing three examples in the training set with ALS input (first column), AI-generation (second column) and ground truth TLS (third column) in each example. Here, we see that our model is able to generate samples at good qualitative approximations of the 3D structural features of TLS distributions conditional on the ALS inputs across the entire vertical tree-stand direction. Figures 6(a) and 6(b) depict instances where the ALS input contains a single tree, which the model accurately approximates, demonstrating its ability to infer complex internal structures from sparse input. Figure 6(c) further illustrates its performance in closely packed tree settings. Here, ground truth TLS shows four trees with tightly packed canopies, while the model generates fewer distinct trunks than the ground truth. Despite this, the output remains visually realistic and spatially consistent with what a human observer might infer from the incomplete ALS scan.
@@
Figure 6: Three examples of ForestGen3D generation at the tree-scale conditioned by ALS inputs in the training set. This qualitative examples showcase the architecture’s capacity to encode and represent the 3D structural details of trees.
Ablation studies to analyze model reconstruction performance where conducted by varying the number of diffusion steps $T\in\{$200, 400, 600, 1000, 2000}\}, diffusion parameters $\beta_{0}=10^{-4}$, $\beta_{T}=0.02$, and batch size 16 constant. These experiments aim to analyze the error of the model at training time and also the spatial containment consistency in eq.(2). We track reconstruction error across training iterations using both Chamfer Distance (CD) and Earth Mover’s Distance (EMD), computed between generated and ground truth TLS point clouds on a held-out validation set of 300 tree examples. As shown in Figs. 7(a)-7(b), both CD and EMD consistently decrease with training iterations across all diffusion schedules, indicating convergence and improving fidelity. Notably, models with higher diffusion steps (e.g., $T=1000$) exhibit slower convergence but ultimately achieve slightly lower reconstruction error, while smaller $T$ values converge quickly with slightly higher error floors.
@@
Figure 7: Chamfer Distance (CD) and Earth’s movers distance (EMD) errors in the validation set as a function of training iterations. Both metrics consistently decrease with training iterations across all diffusion schedules.
@@
(a) Expected Point Containment (EPC)
(b) Chamfer Distance (CD) versus EPC
(c) Earth Movers Distance (EMD) versus EPC
Figure 8: Relationship between containment behavior measured by expected point containment (EPC) and reconstruction error metrics in validation set. There is a strong inverse correlation between EPC and error metrics acrross all denoising time-step schedules $T$: as EPC increases, both Chamfer Distance (CD) and Earth’s movers distance (EMD) rapidly decrease.
@@ 3.3 Testing dataset
For testing, we used 50 TLS scans collected from geographic regions that do not overlap with those of the training and validation areas, though sharing similar ecosystem characteristics and collected with the same sensors used for the training data. We evaluated the model’s performance across three spatial scales. At the tree scale, we followed the same co-registration and tree detection procedures used during training, resulting in 1,457 individual test trees. This setup enabled direct, one-to-one comparisons between generated samples and corresponding TLS ground truth reference data at high structural fidelity. At the plot scale, we constructed test plots by aggregating ALS data over spatial extents corresponding to each TLS scan, truncating ALS coverage to match the TLS field of view (e.g., 25 meters in radius ). This allowed for evaluation of the model’s ability to reconstruct localized forest structure. Finally, at the regional scale, we assessed scalability and extrapolation by deploying the model over multiple 200 meter radius areas containing ALS data. These larger, fully independent areas tested the model’s ability to synthesize realistic vegetation structure across broader, ALS only landscapes. This evaluation strategy enables comprehensive assessment of ForestGen3D’s robustness and flexibility from fine-grained, tree-level detail to broad-scale generation in mixed conifer ecosystems.
@@ 3.4 Tree-Scale ALS-Guided Generation
In Fig. 9 we include three representative qualitative comparisons from the test set to illustrate differences between completion/generation approaches across a range of tree structures. First two columns across all three cases represent ALS input and TLS ground truth, while the remaining columns correspond to the completion/generation resulting from the application of different trained models. In all three examples, PCN [21] tends to oversmooth crowns, recovering coarse silhouettes while missing high-frequency branching detail. The other baseline generative models (3D-GAN [22], latent-GAN [23], and PointFlow [24]) produce shapes that appear closer to the TLS shape, yet are less consistent across examples in general, in comparison to those from our ForestGen3D model. For instance, in the second example, latent-GAN (l) yields a form whose stem does not resemble the straight trunk structure of the TLS reference. Similarly, PointFlow in (f) generates a discontinuous tree stem of larger a diameter compared to ground truth TLS. 3D-GAN and ForestGen3D produce consistently closer results to TLS with canopy structure close to that of ALS and coherent stem-to-crown transitions, recognizable branching patterns, and overall shapes that better match human expectations of realistic tree architecture. These qualitative outcomes mirror the quantitative CD and EMD error, and EPC consistency results below each figure, highlighting the capabilities of generative models to produce realistic structural properties of forests.
@@
Figure 9: Qualitative and Quantitative comparison of ALS-conditioned tree structure generation across three representative examples. (a,h,o) ALS input and (b,i,p) TLS reference; (c,j,q) PCN, (d,k,r) 3D-GAN, (e,l,s) latent-GAN, (f,m,t) PointFlow, and (g,n,u) ForestGen3D (ours). PCN oversmooths crowns, producing coarse silhouettes with little fine-scale detail. The other baselines generate shapes that more closely resemble TLS references but with inconsistencies such as distorted trunks (latent-GAN in (l)) or more discontinuous stems (PointFlow in (f)). 3D-GAN and ForestGen3D yield more coherent reconstructions, with ForestGen3D showing smaller metric errors with respect to ground truth. Pairwise point cloud distance metrics included are Chamfer Distance (CD), Earth’s movers distance (EMD) and Expected Point containment (EPC).
@@
Table 1: Quantitative generation and completion errors evaluated on the test set. Metrics include Chamfer Distance (CD), Earth Mover’s Distance (EMD), and Expected Point Containment (EPC). CD and EMD quantify geometric similarity between generated point clouds and TLS references, while EPC measures spatial containment with the ALS convex envelope when TLS ground truth is unavailable. Lower CD and EMD values and higher EPC indicate better reconstruction quality and structural fidelity.
@@
Table 1 shows a quantitative evaluation summarizing the CD/EMD and EPC metrics over the test dataset of 1457 test examples. This table summarizes the resulting metrics for the completion PCN [21] method and the generative models including 3D-GAN [22], latent l-GAN [23], PointFlow [24] and our ForestGen3D denoising diffusion based model. For reference, example CD/EMD/EPC values are also reported in Fig. 9 for their corresponding examples. Beyond the numerical values, the metrics highlight important differences between the approaches. PCN, while competitive in CD, performs worst in EMD, reflecting its bias toward capturing coarse, low frequency geometry at the expense of fine structural detail. Latent-GAN shows the largest CD, due to plausible over-compressing structural features in the latent layers, while 3D-GAN improves CD, it still remains above our model. PointFlow achieves stronger balance across CD and EMD, though it is also higher than the performance of our model. In contrast, ForestGen3D consistently achieves the lowest CD and EMD, while also attaining the highest EPC, indicating superior alignment with TLS reference structure and better recovery of fine-scale details. These results quantitatively demonstrate that ForestGen3D more effectively learns the underlying distribution of forest structure from ALS/TLS data, achieving overall improved performance compared to both completion/generation-based baselines.
@@ 3.5 Plot-scale ALS-Guided Generation
We also conduct both qualitative and quantitative analyses at the plot scale, centered at TLS scans and covering a 25 meter radius. For the qualitative evaluation, we visually compare structural reconstructions obtained from ALS, TLS, ALS+TLS and a hybrid approach combining ALS with ForestGen3D generation enhancement. For each plot, individual trees were first detected using 3D bounding boxes, and our ForestGen3D model was then applied to generate plausible understory structure beneath each detected tree. Regions lacking tree or shrub like vegetation are left unaltered under the assumption that they do not suffer from occlusion in ALS. To support the quantitative assessment, we perform an evaluation in the downstream task of vegetation biometric estimates derived from each data source: ALS only, TLS only, ALS+TLS and ALS enhanced with ForestGen3D. The combined ALS+TLS data serves as ground truth, given its ability to capture detailed, high-resolution vegetation structure across the entire vertical extent of the tree, including the ground, overstory and understory components. Figure 10 illustrates four representative testset examples (one per row) at plot-scale. The first column illustrates the ALS point cloud, the second column shows the corresponding co-registered ground truth TLS scan, and the third column presents the results of aggregating the ALS scan with the generated under-story from our ForestGen3D model. Note that in all examples, the ALS plots suffer from either missing or highly sparse understory information between the ground and tree foliage, as can be observed in the images in the first column where points are missing. TLS, on the other hand, provides high resolution in the elevation/altitude direction, as shown in the second column, where the main tree stem is finely resolved along the vertical tree-stand direction. The third column, showing the point cloud from ALS+ForestGen3D, illustrates that our approach is capable of enhancing or generating realistic understory structure at the plot scale, where the generated stems and ground features fill in gaps and transition smoothly from the available ALS data.
@@
(f) Ex.4:ALS + ForestGen3D.
(g) Ex.2:ALS.
(h) Ex.2:TLS.
(i) Ex.2:ALS + ForestGen3D.
(j) Ex.3:ALS.
(k) Ex.3:TLS.
(l) Ex.3:ALS + ForestGen3D.
Figure 10: Representative 3D structural comparisons across sensing sources. Each row corresponds to a distinct 25 m radius plot with overlapping ALS and TLS coverage. Columns show ALS (left), TLS (middle), and ALS+ForestGen3D (right). ForestGen3D reconstructs missing vertical stem and mid-story structure absent in ALS, smoothly transitioning from ALS-observed canopy and ground returns into the generated canopy interior and sub-canopy regions.
3.5.1 Do ForestGen3D-Derived Biometrics Approximate TLS-Derived Distributions?
Biometric estimation from 3D point clouds is treated here as a downstream evaluation task rather than the primary objective of ForestGen3D. The goal of this analysis is to assess whether the additional sub-canopy structure generated by the model, while respecting ALS-provided geometry, produces biophysical metric distributions that are consistent with those derived from ALS+TLS. Accordingly, we first evaluate similarity at the level of biometric distributions, which offers a robust and sensor-agnostic assessment of structural realism, and then complement this analysis with per-tree RMSE-based experiments where reliable cross-source tree correspondences are available. This combined evaluation mitigates biases introduced by source-specific tree detection and segmentation differences while still enabling direct quantitative comparison when appropriate.